OpenStudy (anonymous):

Find the value of tan(3pi/16)

OpenStudy (anonymous):

.0102812 if you put it into calculator

OpenStudy (anonymous):

I need the exact value please... How not sure how... :( I know that we need to use the double angle identities.

OpenStudy (freckles):

hi

OpenStudy (anonymous):

Hi freckles! :)

OpenStudy (anonymous):

I need some help...

OpenStudy (freckles):

you know 3pi/4 is on the unit circle right?

OpenStudy (freckles):

and we know 1/2*1/2*3pi/4 is 3pi/16

OpenStudy (freckles):

So it sounds like we are going to have to use the lovely half angle identity we used the other day

OpenStudy (freckles):

\[\tan(\frac{\theta}{2})=\frac{\sin(\theta)}{1+\cos(\theta)}\] have you seen this identity before?

OpenStudy (anonymous):

nope...

OpenStudy (freckles):

\[\tan(u)=\frac{\sin(2 u)}{1+\cos(2u)}\] maybe you seen it written this way but technically this is the same identity I just mentioned just where theta is 2u

OpenStudy (anonymous):

I learned this one: \[\tan2x=\frac{ 2tanx }{ 1-\tan^2x }\]

OpenStudy (freckles):

ok that can also work

OpenStudy (freckles):

replace 2x with u so x is u/2 \[\tan(u)=\frac{2 \tan(\frac{u}{2})}{1-\tan^2(\frac{u}{2})}\] \[u=\frac{3\pi}{4} \\ \tan(\frac{3\pi}{4})=\frac{2 \tan(\frac{ 3 \pi}{8})}{1- \tan^2(\frac{3\pi}{8})}\] we can evaluate tan(3pi/4) using unit circle then we want to try to solve the equation for tan(3pi/8) then we will go about using the half angle identity above to find tan(3pi/16)

OpenStudy (anonymous):

okay

OpenStudy (freckles):

let me know if you want me to do the first round as an example

OpenStudy (freckles):

the second round is basically the same task

OpenStudy (anonymous):

I got this so far: \[\tan^2(\frac{ 3\pi }{ 8 })-2\tan(\frac{ 3\pi }{ 8 })-1=0\]

OpenStudy (freckles):

that is awesome

OpenStudy (freckles):

that is quadratic equation in terms of tan(3pi/8)

OpenStudy (freckles):

you can use the quadratic formula or complete the square to solve q

OpenStudy (anonymous):

I'm not sure how to do that in terms of tan(3pi/8)

OpenStudy (freckles):

pretend v is tan(3pi/8) \[v^2-2v-1=0 \\ v^2-2v=1 \\ v^2-2v+1=1+1 \\ (v-1)^2=2 \\ v-1= \pm \sqrt{2} \\ v=1 \pm \sqrt{2} \\ \text{ but remember } \tan(\frac{3\pi}{8}) \text{ is positive since } 0 < \frac{3\pi}{8} < \frac{\pi}{2} \\ \\ \text{ so } v=1+\sqrt{2} \\ \text{ or since } v=\tan(\frac{3\pi}{8}) \\ \tan(\frac{3\pi}{8})=1+\sqrt{2} \\ \\ \text{ now you need to look at solving this equation for } \tan(\frac{3\pi}{16}) \\ \tan(\frac{3\pi}{8})=\frac{2 \tan(\frac{3\pi}{16})}{1-\tan^2(\frac{3\pi}{16})}\]

OpenStudy (anonymous):

Oh I see, so than it would be \[1+\sqrt2=\frac{ 2\tan(3\pi/16) }{ 1-\tan^2(3\pi/16) }\]

OpenStudy (freckles):

this quadratic will turn out uglier but it is still doable

OpenStudy (anonymous):

I need some help...

OpenStudy (freckles):

\[(1+\sqrt{2})(1)-(1+\sqrt{2})\tan^2(\frac{3\pi}{16})=2 \tan(\frac{3\pi}{16}) \\ -(1+\sqrt{2}) \tan^2(\frac{3\pi}{16})-2 \tan(\frac{3\pi}{16})+(1+\sqrt{2})=0 \\ \text{ \let } x=\tan(\frac{3\pi}{16}) \\ -(1+\sqrt{2})x^2-2 x+(1+\sqrt{2})=0 \] a is that funny thing in front of x^2 b is the thing in front of x the thingy left over, the constant part) is c

OpenStudy (freckles):

I would just plug into the quadratic formula from there

OpenStudy (freckles):

\[x=\frac{-(-2) \pm \sqrt{(-2)^2+4(1+\sqrt{2})(1+\sqrt{2})}}{-2(1+\sqrt{2})}\]

OpenStudy (freckles):

do you have something like this yet?

OpenStudy (freckles):

since again 3pi/16 is between 0 and pi/2 we must choose between the + or the minus in the numerator we want our tan(3pi/16) to be positive and since the bottom is negative we need our top to be negative so we are going to go with the -

OpenStudy (freckles):

you can simplify that answer you have above too

OpenStudy (anonymous):

Sorry for the wait. I got this: \[\tan(3\pi/16)=\frac{ -1+\sqrt{4+2\sqrt{2}} }{ 1+\sqrt2 }\] ??

OpenStudy (freckles):

yep yep

OpenStudy (freckles):

and if you wanted to check it that fraction on the right hand side gives: http://www.wolframalpha.com/input/?i=%28-1%2Bsqrt%284%2B2+sqrt%282%29%29%29%2F%281%2Bsqrt%282%29%29 and tan(3pi/16) gives http://www.wolframalpha.com/input/?i=tan%283pi%2F16%29

OpenStudy (freckles):

both say .6681786 blah blah

OpenStudy (anonymous):

Thanks! @freckles :)